A bad contains marbles of four different colors:

  • 8 blue marbles
  • 24 red marbles
  • 17 green marbles
  • 4 yellow marbles

the probability of picking a blue marble at random out of the bag is


1
Expert's answer
2021-05-18T13:36:32-0400

The probability of picking a blue marble at random out of the bag is "= \dfrac{8}{8+24+17+4} = \dfrac{8}{53}"